OpenStudy (blondisaurus):

Factor by grouping m^2n+mnp-m-p

OpenStudy (anonymous):

m^2n+mnp-m-p = mn (m+p) -1(m+p) = (mn-1) (m+p) are the required factors...

OpenStudy (anonymous):

@Blondisaurus

OpenStudy (blondisaurus):

Still alittle confused but ok. Thanks

OpenStudy (anonymous):

@Blondisaurus Whats the confusion?

OpenStudy (blondisaurus):

Nevermind, I understand now lol. I didnt know you simply split the equation in two and factored out each side...I get it. Thanks again!